bhagwanti inherited Rs.12000/-. she invested part of it as 10% and the rest at 12%. Her annual income from these investment is Rs.1280/- how much did she invest at each rate.
Answer by stanbon(48546)

bhagwanti inherited Rs.12000/-. she invested part of it as 10% and the rest at 12%. Her annual income from these investment is Rs.1280/- how much did she invest at each rate.
-----------------------
Let amount invested at 12% be "x"; Amount of interest = 0.12x Rs
Amount invested at 10% = "12000-x"; Amt of interest = 0.10(12000-x)=1200-0.1x Rs
==========================
EQUATION:
interest + interest = 1280 Rs
0.12x + 1200-0.1x = 1280
0.02x = 80
x = 4000 Rs (amount invested at 12%
12000-x = 8000 Rs ( amount invested at 10%
==================
Cheers,
Stan H.